OpenStudy (anonymous):

Which is an example of the commutative property? A. –13 + 27 = 27 + (–13) B. –21 + (11 + 13) = (–21 + 11) + 13 C. 4(–3 + 5) = 4(–3) + 4 ∙ 5 D. –(–8 + 3) = 8 + (–3)

OpenStudy (anonymous):

Hint: a * b = b * a

OpenStudy (anonymous):

@studyislandhelp

OpenStudy (calculusxy):

Hmm... we know that (a+a)+b=b+(a+a)

OpenStudy (anonymous):

is it D?

OpenStudy (calculusxy):

Now a commutative property is basically like "communicating" to say the truth. Whatever is on one side it is one the other side also. So what do you think is practically the same on both sides?

OpenStudy (anonymous):

B

OpenStudy (calculusxy):

And don't get confused due to parentheses.

OpenStudy (calculusxy):

Affirmatively correct ! :)

OpenStudy (anonymous):

yes!

OpenStudy (calculusxy):

if you like my answer you can be give a medal or be my fan :)

OpenStudy (anonymous):

lol